### 17 Mar 2026 v2

# SWE-CI: Evaluating Agent Capabilities in Maintaining Codebases via Continuous Integration

Abstract
Large language model (LLM)-powered agents have demonstrated strong capabilities in automating software engineering tasks such as static bug fixing, as evidenced by benchmarks like SWE-bench. However, in the real world, the development of mature software is typically predicated on complex requirement changes and long-term feature iterations — a process that static, one-shot repair paradigms fail to capture. To bridge this gap, we propose SWE-CI, the first repository-level benchmark built upon the Continuous Integration loop, aiming to shift the evaluation paradigm for code generation from static, short-term functional correctness toward dynamic, long-term maintainability. The benchmark comprises 100 tasks, each deriving from a real-world code repository with a development history spanning an average of 233 days and 71 consecutive commits. SWE-CI requires agents to systematically resolve these tasks through dozens of rounds of analysis and coding iterations. SWE-CI provides valuable insights into how well agents can sustain code quality throughout long-term evolution.
1 Introduction
Automating software engineering has long been a central objective in the field of artificial intelligence. In recent years, breakthroughs in large language models (LLMs) have lent substantial momentum into this pursuit — from code completion and test generation to end-to-end program repair, LLM-driven agents have demonstrated capabilities rivaling those of human developers across multiple benchmarks. The concurrent evolution of coding benchmarks has played a pivotal role in this progress, providing both rigorous capability measurement and clear research direction.
At the code generation level, HumanEval [3], MBPP [1] and LiveCodeBench [5] established the single-file synthesis paradigm. At the repository level, SWE-bench [6] introduced the “Issue-to-PR” paradigm, requiring models to generate patches within complete repository contexts. At the agent interaction level, Terminal-bench [8] and $\tau$-bench [10] further broadened the evaluation scope to encompass terminal operations and multi-turn tool-use. Together, these efforts have established a multi-granularity and multi-scenario evaluation ecosystem for code intelligence.
Despite the breadth and depth of this ecosystem, its underlying paradigm exhibits one fundamental limitation: Existing benchmarks almost exclusively focus on measuring the agent’s ability to write functionally correct code. However, in the real world, successful software is rarely achieved overnight; it is the result of long-term maintenance. Lehman’s Laws reveal that software quality inherently degrades as maintenance progresses [7], while classic literature has established that maintenance activities account for 60% to 80% of total software lifecycle costs [2]. Therefore, there is an urgent need to design new benchmarks to effectively reflect the ability of models to maintain code.
The reason this capacity has long been absent from evaluation is rooted in the prevailing benchmark paradigm itself. From HumanEval and LiveCodeBench to SWE-bench and Terminal-Bench, existing benchmarks universally adopt a snapshot-style protocol: the agent receives a single, complete requirement and produces a one-shot solution. Under this paradigm, an agent that hard-codes a brittle fix and one that writes clean, extensible code may both pass the same test suite — their difference in maintainability is simply invisible. It becomes visible only when the codebase must evolve: new requirements arrive, interfaces change, and modules must be extended. At that point, the cost of earlier design decisions compounds, and an agent that routinely produces poorly structured code will find each successive modification harder, eventually failing to keep pace. This yields a critical insight: an agent’s ability to maintain code can only be revealed through long-term evolution, where the consequences of past decisions accumulate over successive changes.
Building upon this insight, we propose SWE-CI (SoftWare Engineering – Continuous Integration), a novel benchmark designed to evaluate how well agents maintain code across long-term code evolution. SWE-CI comprises 100 tasks, each defined by a base commit and a target commit from a real-world repository, with an average of 233 days and 71 consecutive commits of authentic evolutionary history in between. SWE-CI employs an Architect–Programmer dual-agent evaluation protocol: starting from the base commit, the agents execute a CI-loop that iteratively generates requirements, modifies source code, and runs tests, with the ultimate objective of passing all tests associated with the target commit. SWE-CI introduces EvoScore (Evolution Score) as a proxy metric: it measures functional correctness on future modifications, so that agents whose earlier decisions facilitate subsequent evolution score higher, while those that accumulate technical debt see progressively declining performance.
We conduct extensive experiments with a total consumption of more than 10 billion tokens. Results reveal that, despite substantial progress on functional correctness, state-of-the-art models still struggle when tasked with sustaining code quality over extended evolution. We further provide comprehensive, fine-grained analyses of the evaluation results, offering valuable insights into the coding capabilities of LLM-based agents and demonstrating the distinctive diagnostic value of SWE-CI.
2 Measuring the Agent’s Ability to Maintain Codebase
2.1 Task formalization
We first establish a unified formalization for agentic coding tasks. Let $t$ denote a single unit test and $\mathcal{T}=\{t_{1},t_{2},\ldots,t_{|\mathcal{T}|}\}$ the set of all tests we are interested in. Let $\mathcal{C}$ denote the space of codebases and $\mathcal{R}$ the space of requirements. We further define two functions:
  • $\mathsf{require}_{\mathcal{T}}:\mathcal{C}\times\mathcal{C}\to\mathcal{R}$, which identifies the functional gap between two codebases with respect to $\mathcal{T}$, and produces a requirement document accordingly;
  • $\mathsf{code}_{\mathcal{T}}:\mathcal{R}\times\mathcal{C}\to\mathcal{C}$, which modifies a codebase according to a given requirement and returns an updated codebase.
With these definitions, we observe that many prevailing coding benchmarks [3, 1, 5, 6, 8, 10] follow the snapshot-based evaluation paradigm illustrated in Figure 1. In this paradigm, the requirement depends only on the base codebase $c_{0}$ and the “golden” codebase $c_{\ast}$, i.e., $r\equiv\mathsf{require}_{\mathcal{T}}(c_{0},c_{\ast})$. In practice, the benchmark maintainers pre-generate $r$ and save it as a prompt, so that benchmark users do not need to regenerate it. However, in this paper, we turn to consider an evolution-based evaluation paradigm. The requirement in this paradigm is derived from the current codebase dynamically: $r_{i}=\mathsf{require}_{\mathcal{T}}(c_{i},c_{*})$, and the codebase is updated accordingly: $c_{i+1}=\mathsf{code}_{\mathcal{T}}(c_{i},r_{i})$. This iterative loop ensures that the consequences of earlier modifications propagate into subsequent iterations, making the agent’s long-term decision quality observable.
Figure 1: Unlike previous benchmarks, SWE-CI proposes an evolution-based evaluation. The red and blue arrows represent the actions of functions $\mathsf{require}$ and $\mathsf{code}$, respectively. Dashed lines indicate processes that are unknown to the user.
2.2 Normalized Change
Most benchmarks treat passing all test cases as the gold standard for functional correctness. However, in software engineering, some features require long-term planning and incremental development rather than a one-shot implementation. Moreover, during evolution it is common for previously passing tests to be inadvertently broken — a phenomenon known as regression. We therefore need a finer-grained metric that reflects the current state of a codebase $c$, rather than a binary pass/fail verdict. To this end, we introduce the normalized change.
Let $n(c)$ be the number of test cases that $c$ passes:
$n(c)=\sum_{t\in\mathcal{T}}\mathbb{I}(t,c),$ (1)
where the indicator function $\mathbb{I}(t,c)$ equals 1 if and only if unit test $t$ passes on codebase $c$, and 0 otherwise. The normalized change is then defined as:
$a(c)=\left\{\begin{aligned} \frac{n(c)-n(c_{0})}{n(c_{\ast})-n(c_{0})}&\,\,\,\,\text{if}\,\,n(c)\geq n(c_{0}),\\ \frac{n(c)-n(c_{0})}{n(c_{0})}&\,\,\,\,\text{else}.\end{aligned}\right$. (2)
When the agent improves upon the base codebase ($n(c)\geq n(c_{0})$), the numerator is normalized by the total gap $n(c_{\ast})-n(c_{0})$, so that $a(c)=1$ if and only if the agent has closed the gap entirely. When the agent regresses below the baseline ($n(c)<n(c_{0})$), the numerator is instead normalized by $n(c_{0})$, so that a(c)=1a(c)=-1 corresponds to breaking every initially passing test. This asymmetric normalization is deliberate: regardless of how large or small $n(c_{0})$ and $n(c_{\ast})$ are, improvements and regressions are always expressed on a unified, comparable scale — i.e., $a(c)\in[-1,1]$.
2.3 EvoScore
The ISO/IEC 25010 standard defines maintainability as the degree to which software can be modified effectively without introducing defects or degrading existing quality [4]. A key implication is that maintainability is not observable at a single snapshot — it manifests only through successive modifications. Guided by this principle, given the sequence of codebases $(c_{1},\ldots,c_{N})$ obtained from $N$ iterations, we aggregate them into a single scalar, EvoScore, via a future-weighted mean:
$e=\frac{\sum_{i=1}^{N}\gamma^{i}\,a(c_{i})}{\sum_{i=1}^{N}\gamma^{i}}$ (3)
In EvoScore, we set $\gamma\geq 1$ so that later iterations receive greater weight. The rationale directly mirrors the ISO definition: a truly maintainable codebase is one that remains easy to modify as evolution progresses. An agent that sacrifices short-term speed for a cleaner, more extensible design will be rewarded over one that rushes to pass early tests but accumulates technical debt that cripples subsequent evolution. When $\gamma=1$, EvoScore reduces to the average normalized change; as $\gamma$ increases, the metric progressively favors long-term stability over immediate gains.
3 SWE-CI
Figure 2: Data curation process of SWE-CI.
3.1 Data curation
As shown in Figure 1, our goal is to obtain a number of base-codebase/oracle-codebase pairs and to let the agent iteratively evolve the former toward the latter, measuring its ability to maintain code throughout this process. Each such pair can be viewed as two chronologically ordered commits within the same repository. Concretely, the construction of SWE-CI is carefully orchestrated as follows:
Step 1: Repository Collection.
Unlike SWE-Bench and similar benchmarks that draw exclusively from a handful of well-known open-source prjects, we cast a wider net by searching across all Python repositories on GitHub. We then apply the following filtering criteria: (1) the repository has been actively maintained for at least three years; (2) it has accumulated more than 500 stars; (3) it contains configuration and dependency files (e.g., pyproject.toml and lockfiles) as well as a suite of unit tests; and (4) it is released under a permissive license such as MIT or Apache-2.0. After applying these filters, 4,923 repositories remain.
Step 2: Commit Span Extraction.
For each surviving repository, we retain only its main branch, reducing the history to a linear sequence of commits. We then compare the dependencies of consecutive commits along this sequence and identify all maximal subsequences within which the dependencies remain unchanged. The two endpoints of every such subsequence naturally form a candidate base/oracle pair. We further discard pairs whose total number of modified lines of code is below 1,000, as such pairs represent insufficient evolutionary distance. This process yields 8,311 candidate pairs.
Step 3: Environment Construction.
For each candidate pair, we automatically generate a Dockerfile based on the configuration and dependencies of the oracle codebase and snapshot the resulting runtime environment. We then execute the oracle codebase’s unit test suite within this environment to verify its correctness. To improve data retention, we introduce a self-repair mechanism: whenever the test suite fails to launch due to a missing dependency, we detect the failure and dynamically inject the required dependency into the Dockerfile to build a new environment. This mechanism substantially increases the number of viable candidate pairs. Pairs whose failures stem from other reasons are discarded. After this step, 1,458 candidate pairs and their runtime environment snapshots remain.
Step 4: Case Filtering.
Finally, we apply three further rounds of filtering to ensure the quality of the final dataset. First, within the runtime environment snapshot constructed in Step 3, we run the oracle codebase’s test suite against the base codebase. Any candidate whose tests fail to launch is discarded. Second, we compare the test reports produced by the base and oracle codebases on the same test suite; candidates for which the difference in the number of passing tests is fewer than five are removed. After these two automated filters, 137 candidates remain. In the last round, we rank the surviving candidates by their time span and number of intervening commits, and select the top 100 to form the final SWE-CI benchmark.
The final SWE-CI benchmark comprises 100 samples drawn from 68 distinct repositories. On average, each base/oracle pair is separated by 233 days and 71 consecutive commits of real-world development history. In every pair, the transition from the base to the oracle codebase involves at least 500 lines of modified source code, excluding changes to test files. Each sample is shipped with the complete source code and a pre-built Docker environment to ensure reproducibility. These statistics confirm that SWE-CI captures substantial, long-term evolutionary episodes rather than trivial incremental changes.
3.2 Dual-agent evaluation protocol
Figure 3: SWE-CI uses an architect-programmer dual-agent workflow to model the continuous integration cycle of professional software teams in the real world.
As described in Figure 1, SWE-CI adopts evolution-based evaluation. To support this setting, we introduce an Architect-Programmer dual-agent protocol. The Architect identifies functional gaps and issues requirements; the Programmer implements them. Their collaboration reproduces the CI loop in real-world development, enabling fine-grained observation of how well agents maintain code.
Architect agent.
Based on the test gap between the current code and the oracle code, the Architect is tasked with producing a high-level requirements document in natural language. We prompt the architect to organize its behavior into three steps: ❶ Summarize. Architect reviews all failing tests, identifies root causes, and identifies source code files that need further inspection; ❷ Locate. Architect examines the source code and attributes failures to concrete deficiencies in the current implementation; ❸ Design. Based on these deficiencies, architect devises an improvement plan and produces the final requirements document. Two writing conventions are further imposed on requirement document. ➀ Incremental. The document should contain no more than five of the most urgent requirements, avoiding the pitfall of over-designing in a single iteration. ➁ High-level. the requirements should focus on describing expected behavior of code using neural language, leaving concrete implementation choices to the programmer. The core purpose of these specifications is to ensure that requirement documents meet the needs of real-world continuous integration processes.
Programmer.
The programmer’s responsibility is to maintain the code according to the requirements document. Programmer behavior is also standardized into three steps: ❶ Comprehend. Programmers understand high-level language requirements and translate them into explicit code specifications. ❷ Plan. Programmers plan the programming effort required to implement these specifications. ❸ Code. Programmers put these plans into practice and try to fulfill the requirements.
In this protocol, the Programmer is driven by the requirements document rather than directly by the test gap — a deliberate design choice that aligns with the rapid iteration philosophy of continuous integration. To this end, the Architect is required to distill the most pressing requirements from the full set of failures, allowing the Programmer to focus on fast, targeted development without being overwhelmed by the full scope of changes.
4 Experiments
4.1 Experiment setting
We use pytest and pytest-json-report as the testing framework, with a timeout of 3600 seconds per test run. iFlow CLI [9] serves as the default agent framework, and the maximum number of iterations in the dual-agent evaluation protocol is set to 20. Unless otherwise specified, the Architect Agent and the Programmer Agent share the same underlying base model.
4.2 Results
Observation 1: The code maintenance capabilities of LLMs are advancing at an accelerating pace (Figure 4).
Our extensive evaluation of 18 models from 8 different providers reveals a consistent pattern: within the same provider family, newer models always achieve higher scores, with models released after 2026 showing markedly larger gains than their predecessors. This suggests that the code capabilities of current LLMs are rapidly evolving beyond static bug-fixing toward sustained, long-term code maintenance. Among all evaluated models, the Claude Opus series demonstrates a commanding lead throughout the entire observation period, with GLM-5 also standing out as a strong performer.
Figure 4: The EvoScore variation of state-of-the-art models from 8 providers on SWE-CI.
Observation 2: Different provider place varying degrees of emphasis on code maintainability. (Figure 5).
We vary the value of $\gamma$ to examine how model rankings shift accordingly. When $\gamma<1$, EvoScore assigns higher weights to earlier iterations, favoring models that prioritize immediate gains from code modification. Conversely, when $\gamma>1$, later iterations are rewarded, giving an advantage to models that optimize for long-term improvement (i.e., prioritize code maintainability). We find that preferences vary considerably across providers, while models within the same provider tend to exhibit consistent tendencies. Specifically, MiniMax, DeepSeek, and GPT all show a preference for long-term gains, whereas Kimi and GLM lean toward short-term returns. Qwen, Doubao, and Claude, by contrast, remain relatively stable across different settings. We conjecture that this reflects differences in training strategies adopted by different providers, while the relative consistency within each provider suggests that their internal training pipelines remain largely stable.
Figure 5: The model’s EvoScore ranking changes as $\gamma$ increases. When $\gamma>1$, higher-ranking models indicate better codebase maintenance.
Observation 3: Current LLMs still fall short in controlling regressions during long-term code maintenance. (Figure 6).
Regression is a core metric for measuring software quality stability — if a unit test passes before a code change but fails afterward, the change is considered to have introduced a regression. In software maintenance, regressions must be strictly monitored. Once a regression occurs, it not only directly impacts user experience, but can also lead to systematic quality degradation as the number of changes accumulates over long-term maintenance. To this end, we measure in SWE-CI the proportion of samples in which no regression occurs throughout the entire code maintenance process, referred to as the zero-regression rate, to evaluate the stability of models in continuous maintenance scenarios. Experimental results show that most models achieve a zero-regression rate below 0.25, with only two models in the Claude-opus series exceeding 0.5, indicating that current LLMs still struggle to reliably avoid regressions in long-term code maintenance. This suggests that, although LLMs have shown significant improvements in snapshot-based code modification tasks, they still face substantial challenges in fully automated, long-term, and multi-round software development and maintenance scenarios.
Figure 6: All models are sorted from smallest to largest by the zero regression rate.
Acknowledgements
The authors thank colleagues from Alibaba Group, including Jianan Ye, Cecilia Wang, Yijie Hu, Zongwen Shen, and Mingze Li, for their valuable suggestions and feedback on this paper.
Appendix A Prompts
System Prompt for Architect Agent System Prompt for Programmer Agent